# Global survey of star clusters in the Milky Way VIII. Cluster formation and evolution

### v1

@@ 4.1 Error model
In order to compute a modelled CAMF that includes smoothing by age and mass errors, we need to model the measurement errors of cluster age and mass. We do this using functions $\delta_{\tau}(\mu,\tau)$ and $\delta_{\mu}(\mu,\tau)$ as introduced in Sect. 3.
As the age errors we used are relatively small and mostly the same across all ages, we use simply
@@
as this is consistent with the uncertainties claimed by Kharchenko et al. (2013).
For the mass errors, we considered correlations of the logarithmic cluster age and mass with the logarithmic mass error and found that the error is largely independent of the mass, but does depend on age.
@@
When using this error model to compute an error-smoothed model CAMF following Eq. (14) to compare to the observed error-smoothed CAMF constructed according to Eq. (7), we apply a factor of $\sqrt{2}$ to the modelled errors to account for the fact that in addition to the explicit application of the uncertainties to the CAMF, the observed values of cluster ages and masses are already scattered from their true values on a scale estimated by their given uncertainties.
@@ 4.3 Kullback-Leibler divergence
Our second approach to judge the correspondence of the model to the observed CAMF was based on the Kullback-Leibler divergence (KLD).
For two discrete probability distributions $p,\,q$, the KLD
@@
where $\bar{\sigma}=\frac{1}{N_{\mathrm{bin}}}\sum\limits_{k,l}^{N_{\mathrm{bin}}}\hat{\sigma}_{k,l}$, for the KLD between observed and modelled CAMFs including measurement uncertainties. As with the $\chi^{2}$-statistic, we include an artificial constant offset regulated by $\epsilon$ which avoids singularities that arise when the modelled CAMF goes to zero.
@@ 4.4 Expected likelihood
with $\Sigma=\sum\limits_{k=1}^{N_{\mathrm{Cl}}}S_{k}^{-1}$, represents the fraction of clusters that have the same properties as the $i$-th cluster of our sample, using the same argument by which the binned normalised CAMF represents a discrete probability distribution of cluster age-mass intervals.
We can use this expected log-likelihood as a measure of model fitness by the following argument: Given a representative sample $\mathcal{S}$ of $N$ clusters and a prior probability of the model $P_{\mathrm{prior}}$, the model likelihood given the sample is, by Bayes’ theorem,
@@
where $P(\mathcal{S})$ is the (unknown) probability of the sample, which does not depend on the choice of the model.
Since for $N\rightarrow\infty$
$\frac{1}{N}\sum\limits_{i=1}^{N}\ln P_{\mathrm{O}}(m_{i},t_{i})\rightarrow E[\ln P_{\mathrm{O}}(m_{i},t_{i})],$ (39)
@@
The reason that we do not use the likelihood of the observed cluster sample directly lies in the fact that our sample features individual completeness limits. As in Eq. (36), we can construct relative weights $n_{i}$ for the individual cluster log-likelihoods using the completeness areas, however it is not readily evident what effective sample size $N_{\mathrm{eff}}=\sum\limits_{i=1}^{N_{\mathrm{Cl}}}n_{i}$ should be used to construct the total sample likelihood. As such, we use $\sum\limits_{i=1}^{N_{\mathrm{Cl}}}n_{i}=1$ to estimate only the expected log-likelihood for a single cluster.
@@ 4.5 Parameter constraints and priors
From the list of parameters as given in Table 1, we vary the parameter vector
@@
which gives us a total of $N_{\mathrm{par}}=11$ free parameters.
In particular, since we fit model to data in the logarithmic age-mass plane, we vary parameters that represent times and masses such as $M_{\ast},\,m_{\mathrm{S}},\,M_{\mathrm{br}}$ and $t_{\mathrm{V}}$ in logarithmic space directly, likewise $c$, which has dimension of inverse time, and $n_{\mathrm{b}}$, which is a mass fraction.
Regarding the remaining parameters, $k$ normalises the CIMF and is thus fixed by the other CIMF parameters. Since $k$ has inverse mass dimension, we will in the context of parameter fitting also give its value in logarithmic space.
The parameter $\beta$ is fixed given all other model parameters by matching the modelled total cluster surface density with the observed one (see for this Sect. 3.1).
The parameters of $\mu_{\mathrm{L}}(t)$ remain fixed to the values from Lamers et al. (2010) for $Z=0.008$ given in Table 1. This is because the mass loss from stellar evolution gives a comparatively small contribution, such that varying it gives little benefit. Moreover, $\mu_{\mathrm{L}}(t)$ depends on underlying physical assumptions regarding stellar evolution and the stellar IMF that are not easily translated into constraints for the parameters $p_{0},\,p_{1},\,p_{2}$ and $p_{3}$. As such, in the context of parameter fitting, one approach would be to instead vary the underlying parameters such as the metallicity and compute the corresponding values of the $p_{i}$ by interpolating from tabulated values.
@@
The remaining parameters are constrained in consideration of three aspects: The definition of the model functions, the computability of the model CAMF and the physicality of the model. The first aspect refers to the fact that certain parameters are required to have certain properties, such as, for instance, non-negativity, by the definition of their corresponding model functions. As such, we require non-negativity for $x_{1},s,a_{2}$, as well as $x_{1}<x_{2}$, $a_{1}<1$ and $\log n_{\mathrm{b}}<0$. The second aspect mainly refers to the requirement for the lifetime-mass relation to be monotonic such that $m(M,t)$ can be inverted to compute $M$ for given $m,t$. We require $0.11-a_{1}-a_{2}>$ 0.1 instead of simple positiveness in order to avoid poor numerical conditioning. Similarly, we impose finite bounds for $s$ in order to avoid undesirable behaviour of the CIMF in the limits $s\rightarrow 0$ and ${(x_{2}-x_{1})}/{s}\rightarrow 0$. The third aspect refers to the observational and physical constraints we impose on the model from prior knowledge. We know that some clusters reach ages $\tau>9.5$, but others dissolve within the observed range of ages, which gives us constraints for the cluster lifetimes. The existence of clusters of a given mass places constraints on the Schechter cut-off mass. Prior work on N-body dynamics gives limits on the timeframe during which violent relaxation will take place (see Shukirgaliyev et al. 2017, 2018, 2019, 2021).
@@
In Table 2, the ranges to which each parameter is confined are given in the third column, and the more complex constraints involving multiple parameters are listed in the fourth column.
@@
In the second column of Table 2, Gaussian priors with mean $\mu_{\mathrm{G}}$ and standard deviation $\sigma_{\mathrm{G}}$ are given in the form $\mu_{\mathrm{G}}\pm\sigma_{\mathrm{G}}$ for some parameters. These priors represent information that can easily be extracted from the data without performing a full fit in eleven-dimensional parameter space.
@@
As discussed in Just et al. (2023) (see in particular Table 3 therein), the high-mass power-law slope of the CMF is close to $x_{2}=1$ independent of age, while the low-mass slope changes significantly as the cluster population evolves, being shallow initially and steeper for higher cluster ages. Extrapolating from this, we adopt priors for $x_{1}$ and $x_{2}$ that reflect a high-mass slope close to one and a more uncertain, but likely shallow low-mass slope.
@@
Regarding the parameters of violent relaxation, we consider the first two moments of the cluster mass as a function of age, in particular the mean and the standard deviation of the cluster mass. Within the context of our model, the bound mass fraction $n_{\mathrm{V}}(t)$ applies as a factor equally to both of these moments. Assuming that violent relaxation dominates the early mass evolution, we can thus extract information on $n_{\mathrm{b}}$ and $t_{\mathrm{V}}$ from them.
@@
We identify a transition in the cluster masses consistent with our expectations of violent relaxation in the age range of $\tau=7$–7.5. As such, we compare masses for age ranges $\tau=6.5$–7 and $\tau=7.5$–7.8 to estimate $n_{\mathrm{b}}$. We find $\log n_{\mathrm{b}}=-0.42\pm 0.13$ from the mean cluster mass and $\log n_{\mathrm{b}}=-0.53\pm 0.07$ from the standard deviation. These figures, however, likely underestimate the effect of violent relaxation as they neglect mass loss during the first 10 Myr of cluster evolution. The prior we adopt for the bound mass fraction is thus slightly lower. The transition in the mass moments at ages $\tau=7$–7.5 informs the prior for the relaxation timescale $t_{\mathrm{V}}$.
@@
Finally, we consider again the CMF for Initial-Age clusters ($\tau=6.4$–7.3 as per Just et al. (2023)). Here the transition between power-law slopes occurs around $\mu=2.3$. As this age group overlaps with the period of violent relaxation, we can infer that probably $\log M_{\ast}/\mathrm{M}_{\odot}>2.3$ and $\log M_{\ast}/\mathrm{M}_{\odot}+\log n_{\mathrm{b}}<2.3$. As such, the prior for $n_{\mathrm{b}}$ informs our prior for $M_{\ast}$.
@@
For the remaining parameters, we do not find Gaussian priors and instead use flat priors which are uniform within the parameter constraints.
@@ 4.6 Markov Chain Monte Carlo sampling
The parameter fitting is then done using the standard Bayesian approach where both the likelihood $L(\boldsymbol{\theta})=P(\mathrm{sample}|\boldsymbol{\theta})$ of observing our cluster sample given a model with parameter vector $\boldsymbol{\theta}$ and the prior probability $P(\boldsymbol{\theta})$ of the parameters are combined into the posterior probability $\mathcal{P}(\boldsymbol{\theta})$ according to Bayes’ theorem. Logarithmically, we can write
@@
for some normalising constants $\mathcal{P}_{0},\,L_{0}$, and $P_{0}$ for posterior, likelihood, and prior, respectively. Explicitly knowing these normalising constants is not necessary to sample the posterior probability distribution, which is done using a MCMC approach. By constructing a sufficiently large and representative sample of the posterior, we can estimate parameter uncertainties and correlations as well as find best fit values of the parameters by finding parameter vectors which maximise the posterior probability.
In detail, this approach is varied here in how the likelihood $L(\boldsymbol{\theta})$ is estimated. For the $\chi^{2}$-statistic (see Eq. (28)), we use
@@
where $\chi^{2}_{\mathrm{red}}=\frac{\chi^{2}}{n_{\mathrm{dof}}}$ is the reduced $\chi^{2}$-statistic for $n_{\mathrm{dof}}=n_{\mathrm{bins}}-n_{\mathrm{pars}}-1$ degrees of freedom. We use an offset with $\epsilon=0.005$.
For the KLD (see Eq. (31)), we use
@@
Similarly, we use
$\ln\frac{L}{L_{0}}=\kappa_{\mathrm{EL}}E[\ln P(m_{i},t_{i}|\boldsymbol{\theta})]$ (43)
@@
Both scaling constants were chosen such that the likelihood varies across similar-sized ranges for all three fit statistics.
@@
We use the Python emcee module (Foreman-Mackey et al. 2013) for MCMC sampling, which uses an implementation of the affine-invariant ensemble sampler described by Goodman & Weare (2010). This ensemble sampler runs a number $N_{\mathrm{walker}}>N_{\mathrm{par}}$ of samplers in parallel, using the space spanned by them to sample the high-dimensional parameter space more efficiently. In terms of run setup, we broadly follow the advice of Foreman-Mackey et al. (2013) as well as the emcee documentation. We use $N_{\mathrm{walker}}=52$ for our 11-dimensional parameter space, running the chain in 13 parallel CPU processes using the emcee module’s multiprocessing support. Convergence of the chain is tested using the estimated integrated autocorrelation time $\hat{t}_{\boldsymbol{\theta\theta}}$ of the parameters, as computed using the emcee module. It is further used in taking a more representative set of samples from the final chain by discarding the first $\max(2\hat{t}_{\boldsymbol{\theta\theta}})$ samples as burn-in and taking only every $\min(\hat{t}_{\boldsymbol{\theta\theta}}/2)$-th sample of the remaining chain to ensure independence of the samples. The resulting set of parameter combinations is then taken to be a sufficiently large and representative sample of the posterior.
@@ 5.1 Best-fitting models
We note that while the best-fitting models are those which best agree with the data within each MCMC sample, they are not best-fit models in the sense of representing a local optimum of a fit statistic. In the same sense, the percentile ranges do not constitute a fitting error, but a range of values the parameters typically take for models that agree reasonably well with the data. Nonetheless, for brevity we refer to the MCMC runs in the following as $\chi^{2}$-fit, $D_{\mathrm{KL}}$-fit, and $E[\ln P]$-fit, and to the respective best-fitting models as the $\chi^{2}$-fit model, the $D_{\mathrm{KL}}$-fit model, and the $E[\ln P]$-fit model.
In principle, it is possible to use the best-fitting parameter sets as starting points to find such local optima, but as we will discuss in the following, such a best-fit model may not actually be a useful improvement.
@@
Other features, such as the high-density region of clusters with intermediate ages and masses as well as the low-mass limit are more consistently reproduced across the models considered.
We notice also that our models all produce an overabundance of low-mass clusters ($\mu<1$) at intermediate ages $\tau=7.7$–8.5.

### v2

@@ 4.1 Error model
In order to compute a modelled CAMF that includes smoothing by age and mass errors, we need to model the measurement errors of cluster age and mass. We did this using functions $\delta_{\tau}(\mu,\tau)$ and $\delta_{\mu}(\mu,\tau)$ as introduced in Sect. 3. As the age errors we used are relatively small and mostly the same across all ages, we used simply
@@
as this is consistent with the uncertainties claimed by Kharchenko et al. (2013). For the mass errors, we considered correlations of the logarithmic cluster age and mass with the logarithmic mass error and found that the error is largely independent of the mass, but does depend on age.
@@
When using this error model to compute an error-smoothed model CAMF following Eq. (14) to compare to the observed error-smoothed CAMF constructed according to Eq. (7), we applied a factor of $\sqrt{2}$ to the modelled errors. This was to account for the fact that in addition to the explicit application of the uncertainties to the CAMF, the observed values of cluster ages and masses are already scattered from their true values on a scale estimated by their given uncertainties.
@@ 4.3 Kullback-Leibler divergence
Our second approach to judge the correspondence of the model to the observed CAMF was based on the Kullback-Leibler divergence (KLD). For two discrete probability distributions $p,\,q$, the KLD
@@
where $\bar{\sigma}=\frac{1}{N_{\mathrm{bin}}}\sum\limits_{k,l}^{N_{\mathrm{bin}}}\hat{\sigma}_{k,l}$, for the KLD between observed and modelled CAMFs including measurement uncertainties. As with the $\chi^{2}$-statistic, we included an artificial constant offset regulated by $\epsilon$ which avoids singularities that arise when the modelled CAMF goes to zero.
@@ 4.4 Expected likelihood
with $\Sigma=\sum\limits_{k=1}^{N_{\mathrm{Cl}}}S_{k}^{-1}$. $n_{i}$ represents the fraction of clusters that have the same properties as the $i$-th cluster of our sample, using the same argument by which the binned normalised CAMF represents a discrete probability distribution of cluster age-mass intervals.
We used this expected log-likelihood as a measure of model fitness by the following argument: Given a representative sample $\mathcal{S}$ of $N$ clusters and a prior probability of the model $P_{\mathrm{prior}}$, the model likelihood given the sample is, by Bayes’ theorem,
@@
where $P(\mathcal{S})$ is the (unknown) probability of the sample, which does not depend on the choice of the model. Since for $N\rightarrow\infty$
$\frac{1}{N}\sum\limits_{i=1}^{N}\ln P_{\mathrm{O}}(m_{i},t_{i})\rightarrow E[\ln P_{\mathrm{O}}(m,t)],$ (39)
@@
The reason that we did not use the likelihood of the observed cluster sample directly lies in the fact that our sample features individual completeness limits. As in Eq. (36), we can construct relative weights $n_{i}$ for the individual cluster log-likelihoods using the completeness areas, however it is not readily evident what effective sample size $N_{\mathrm{eff}}=\sum\limits_{i=1}^{N_{\mathrm{Cl}}}n_{i}$ should be used to construct the total sample likelihood. As such, we used $\sum\limits_{i=1}^{N_{\mathrm{Cl}}}n_{i}=1$ to estimate only the expected log-likelihood for a single cluster.
@@ 4.5 Parameter constraints and priors
From the list of parameters as given in Table 1, we varied the parameter vector
@@
which gave us a total of $N_{\mathrm{par}}=11$ free parameters. In particular, since we fit model to data in the logarithmic age-mass plane, we varied parameters that represent times and masses such as $M_{\ast},\,m_{\mathrm{S}},\,M_{\mathrm{br}}$ and $t_{\mathrm{V}}$ in logarithmic space directly, likewise $c$, which has dimension of inverse time, and $n_{\mathrm{b}}$, which is a mass fraction.
Regarding the remaining parameters, $k$ normalises the CIMF and was thus fixed by the other CIMF parameters. Since $k$ has inverse mass dimension, we give in the context of parameter fitting also its value in logarithmic space. The parameter $\beta$ was fixed given all other model parameters by matching the modelled total cluster surface density with the observed one (see for this Sect. 3.1). The parameters of $\mu_{\mathrm{L}}(t)$ remained fixed to the values from Lamers et al. (2010) for $Z=0.008$ given in Table 1. This was because the mass loss from stellar evolution gives a comparatively small contribution, such that varying it gives little benefit. Moreover, $\mu_{\mathrm{L}}(t)$ depends on underlying physical assumptions regarding stellar evolution and the stellar IMF that are not easily translated into constraints for the parameters $p_{0},\,p_{1},\,p_{2}$ and $p_{3}$. As such, in the context of parameter fitting, one approach would be to instead vary the underlying parameters such as the metallicity and compute the corresponding values of the $p_{i}$ by interpolating from tabulated values.
@@
The remaining parameters were constrained in consideration of three aspects: The definition of the model functions, the computability of the model CAMF and the physicality of the model. The first aspect refers to the fact that certain parameters are required to have certain properties, such as, for instance, non-negativity, by the definition of their corresponding model functions. As such, we required non-negativity for $x_{1},s,$ and $a_{2}$, as well as $x_{1}<x_{2}$, $a_{1}<1$ and $\log n_{\mathrm{b}}<0$. The second aspect mainly refers to the requirement for the lifetime-mass relation to be monotonic such that $m(M,t)$ can be inverted to compute $M$ for given $m,t$. We required $0.11-a_{1}-a_{2}>$ 0.1 instead of simple positiveness in order to avoid poor numerical conditioning. Similarly, we imposed finite bounds for $s$ in order to avoid undesirable behaviour of the CIMF in the limits $s\rightarrow 0$ and ${(x_{2}-x_{1})}/{s}\rightarrow 0$. The third aspect refers to the observational and physical constraints we impose on the model from prior knowledge. We know that some clusters reach ages $\tau>9.5$, but others dissolve within the observed range of ages, which gives us constraints for the cluster lifetimes. The existence of clusters of a given mass places constraints on the Schechter cut-off mass. Prior work on N-body dynamics gives limits on the timeframe during which violent relaxation will take place (see Shukirgaliyev et al. 2017, 2018, 2019, 2021).
@@
In Table 2, the ranges to which each parameter was confined are given in the third column, and the more complex constraints involving multiple parameters are listed in the fourth column.
@@
In the second column of Table 2, Gaussian priors with mean $\mu_{\mathrm{G}}$ and standard deviation $\sigma_{\mathrm{G}}$ are given in the form $\mu_{\mathrm{G}}\pm\sigma_{\mathrm{G}}$ for some parameters. These priors represent information that could easily be extracted from the data without performing a full fit in eleven-dimensional parameter space.
@@
As discussed in Just et al. (2023) (see in particular Table 3 therein), the high-mass power-law slope of the CMF is close to $x_{2}=1$ independent of age, while the low-mass slope changes significantly as the cluster population evolves, being shallow initially and steeper for higher cluster ages. Extrapolating from this, we adopted priors for $x_{1}$ and $x_{2}$ that reflect a high-mass slope close to one and a more uncertain, but likely shallow low-mass slope.
@@
Regarding the parameters of violent relaxation, we considered the first two moments of the cluster mass as a function of age, in particular the mean and the standard deviation of the cluster mass. Within the context of our model, the bound mass fraction $n_{\mathrm{V}}(t)$ applied as a factor equally to both of these moments. Assuming that violent relaxation dominates the early mass evolution, we could thus extract information on $n_{\mathrm{b}}$ and $t_{\mathrm{V}}$ from them.
@@
We identified a transition in the cluster masses consistent with our expectations of violent relaxation in the age range of $\tau=7$–7.5. As such, we compared masses for age ranges $\tau=6.5$–7 and $\tau=7.5$–7.8 to estimate $n_{\mathrm{b}}$. We found $\log n_{\mathrm{b}}=-0.42\pm 0.13$ from the mean cluster mass and $\log n_{\mathrm{b}}=-0.53\pm 0.07$ from the standard deviation. These figures, however, likely underestimate the effect of violent relaxation as they neglect mass loss during the first 10 Myr of cluster evolution. The prior we adopted for the bound mass fraction is thus slightly lower. The transition in the mass moments at ages $\tau=7$–7.5 informed the prior for the relaxation timescale $t_{\mathrm{V}}$.
@@
Finally, we considered again the CMF for Initial-Age clusters ($\tau=6.4$–7.3 as per Just et al. (2023)). Here the transition between power-law slopes occurs around $\mu=2.3$. As this age group overlaps with the period of violent relaxation, we can infer that probably $\log M_{\ast}/\mathrm{M}_{\odot}>2.3$ and $\log M_{\ast}/\mathrm{M}_{\odot}+\log n_{\mathrm{b}}<2.3$. As such, the prior for $n_{\mathrm{b}}$ informed our prior for $M_{\ast}$.
@@
For the remaining parameters, we did not find Gaussian priors. Instead, we used flat priors which are uniform within the parameter constraints.
@@ 4.6 Markov chain Monte Carlo sampling
The parameter fitting was then done using the standard Bayesian approach where both the likelihood $L(\boldsymbol{\theta})=P(\mathrm{sample}|\boldsymbol{\theta})$ of observing our cluster sample given a model with parameter vector $\boldsymbol{\theta}$ and the prior probability $P(\boldsymbol{\theta})$ of the parameters are combined into the posterior probability $\mathcal{P}(\boldsymbol{\theta})$ according to Bayes’ theorem. Logarithmically, we can write
@@
for some normalising constants $\mathcal{P}_{0},\,L_{0}$, and $P_{0}$ for posterior, likelihood, and prior, respectively. Explicitly knowing these normalising constants is not necessary to sample the posterior probability distribution, which was done using a MCMC approach. By constructing a sufficiently large and representative sample of the posterior, we can estimate parameter uncertainties and correlations as well as find best fit values of the parameters by finding parameter vectors which maximise the posterior probability.
In detail, this approach was varied here in how the likelihood $L(\boldsymbol{\theta})$ is estimated. For the $\chi^{2}$-statistic (see Eq. (28)), we used
@@
where $\chi^{2}_{\mathrm{red}}=\frac{\chi^{2}}{n_{\mathrm{dof}}}$ is the reduced $\chi^{2}$-statistic for $n_{\mathrm{dof}}=n_{\mathrm{bins}}-n_{\mathrm{pars}}-1$ degrees of freedom. We used an offset with $\epsilon=0.005$.
For the KLD (see Eq. (31)), we used
@@
Similarly, we used
$\ln\frac{L}{L_{0}}=\kappa_{\mathrm{EL}}E[\ln P(m,t|\boldsymbol{\theta})]$ (43)
@@
Both scaling constants were chosen such that the likelihood varied across similar-sized ranges for all three fit statistics.
@@
We used the Python emcee module (Foreman-Mackey et al. 2013) for MCMC sampling, which uses an implementation of the affine-invariant ensemble sampler described by Goodman & Weare (2010). This ensemble sampler runs a number $N_{\mathrm{walker}}>N_{\mathrm{par}}$ of samplers in parallel, using the space spanned by them to sample the high-dimensional parameter space more efficiently. In terms of run setup, we broadly followed the advice of Foreman-Mackey et al. (2013) as well as the emcee documentation. We used $N_{\mathrm{walker}}=52$ for our 11-dimensional parameter space, running the chain in 13 parallel CPU processes using the emcee module’s multiprocessing support. Convergence of the chain was tested using the estimated integrated autocorrelation time $\hat{t}_{\boldsymbol{\theta\theta}}$ of the parameters, as computed using the emcee module. It was further used in taking a more representative set of samples from the final chain by discarding the first $\max(2\hat{t}_{\boldsymbol{\theta\theta}})$ samples as burn-in and taking only every $\min(\hat{t}_{\boldsymbol{\theta\theta}}/2)$-th sample of the remaining chain to ensure independence of the samples. The resulting set of parameter combinations was then taken to be a sufficiently large and representative sample of the posterior.
@@ 5.1 Best-fitting models
We note that while the best-fitting models are those which best agree with the data within each MCMC sample, they are not best-fit models in the sense of representing a local optimum of a fit statistic. In the same sense, the percentile ranges do not constitute a fitting error, but a range of values the parameters typically take for models that agree reasonably well with the data. Nonetheless, for brevity we refer to the MCMC runs in the following as $\chi^{2}$-fit, $D_{\mathrm{KL}}$-fit, and $E[\ln P]$-fit, and to the respective best-fitting models as the $\chi^{2}$-fit model, the $D_{\mathrm{KL}}$-fit model, and the $E[\ln P]$-fit model. In principle, it is possible to use the best-fitting parameter sets as starting points to find such local optima, but as we discuss in the following, such a best-fit model may not actually be a useful improvement.
@@
Other features, such as the high-density region of clusters with intermediate ages and masses as well as the low-mass limit are more consistently reproduced across the models considered. We notice also that our models all produce an overabundance of low-mass clusters ($\mu<1$) at intermediate ages $\tau=7.7$–8.5.